# On the Effectiveness of Code Representation in Deep Learning-Based Automated Patch Correctness Assessment

### 08 Mar 2026 v1

@@ 5.1. Impact of Node Embedding
Objective. As discussed in Section 4.1, tree-based and graph-based representations have demonstrated impressive performance compared with other representations and state-of-the-art APCA approaches. The tree-based and graph-based representations are composed of nodes that represent the code element and edges that represent relationships between different nodes. In particular, each node in the tree and graph is composed of two parts, the node type and the node textual information. The former attempts to distinguish the node from the others, and the latter is a small fragment code snippet from the original program (the first line in each node). In our experiment, we utilize both the node type and the textual information to obtain the embedding to represent each node by default. However, it is unclear whether they both have a positive effect on the prediction results.
@@
Summary: Textual information yields better results than type information due to the rich code semantics for predicting patch correctness.
@@ 5.2. Potential of Pre-trained Code Models
Recently, pre-trained code models (PCMs) have achieved impressive results in SE tasks, including program repair and patch validation. The advanced performance of these models stems from their architecture and parameter scale, while this paper aims to investigate the impact of code representations. Therefore, to eliminate the influence of more advanced models, we did not use PCMs in the main experiments. In this section, we will explore the performance of these PCMs in the context of patch validation. Particularly, we consider an encoder-only PCM CodeBERT and an encoder-decoder PCM CodeT5,
@@ 5.3. Threats to Validity
The first threat to validity is the selection of baselines. In Section 4.2, we consider four APCA techniques as baselines to evaluate the effectiveness of four types of representations. We may fail to consider other techniques (mentioned in Section 6) due to the rapidly growing community and the constant emergence of APCA approaches. However, our selected baselines represent state-of-the-art (e.g., the traditional dynamic-based PATCH-SIM) and are widely-adopted in almost all of previous studies (Lin et al., 2022; Tian et al., 2020). Besides, these selected baselines cover various categories, such as ODS (Ye et al., 2022) with hand-crafted features, Tian et al. (Tian et al., 2020) with code sequences, and CACHE (Lin et al., 2022) with ASTs. Furthermore, our work mainly focuses on empirical evaluations of the effectiveness of different representations. Thus, we believe that the baselines have little impact on results of our work.
@@
The second threat to validity is the evaluation benchmark. In our experiment, we focus on the most popular Defects4J in the APR and APCA literature involving 835 bugs from 17 open-source projects. It is unclear the degree to which the results in our work can be generalized to other benchmarks. Despite other datasets available (Lin et al., 2022), they are quite easy to identify due to simple bug types, such as 99.9% AUC (Zhang et al., 2024a), making it difficult to assess true capabilities. To address this threat, we collect 2,274 patches generated by 39 existing APR tools, to the best of our knowledge, which is the largest patch set of Defects4J in the literature. It should be noted that, with the continuous release of newer APR tools on Defects4J, we observed that their contributions of unique patches (i.e., those not already generated by earlier tools) have become increasingly limited. In fact, some of their patches overlap with patches already included in our dataset. For this reason, while Table 1 does not explicitly list these tools, their outcomes are partially covered, and our dataset remains both comprehensive and the largest available. Besides, we mitigate the potential bias by using multiple evaluation metrics to exhaustively assess the involved techniques. In the future, we plan to continuously update the dataset by incorporating newer APR tools, such as Tare (Zhu et al., 2023) and Recoder (Zhu et al., 2021), to further strengthen the completeness of our benchmark.
@@
The third threat to validity lies in the evaluation setting. First, following most previous studies (Lin et al., 2022; Tian et al., 2020), we evaluate the performance of APCA approaches within a cross-validation setting, which is the common practice in the APCA community. It is possible that the patches generated by different APR tools for the same bug are split into the training set and testing set. However, in a real-world scenario, it would be impossible to use the label from one patch to predict the other for the same bug. Second, there exist more potential code representations and models that are excluded in our experiments due to page limit. Third, we focus on filtering overfitting patches following most prior work and the results may not be generalized to correct patch identification. To address this threat, we construct a large benchmark with 2,274 patches from 39 APR tools, consider 15 representations and 11 models, and select seven baselines from different categories. Besides, considering the scale of our empirical evaluation involving nearly 600 trained models, i.e., (45 models in RQ1 + one model in RQ2 + 22 models in RQ3 + 51 models in Section 5.1) $\times$ 5-fold cross-validation, the current results are enough to demonstrate the performance of code representations in patch correctness prediction. In the future, researchers could further explore the experiments under more potential scenarios, such as the cross-project evaluation and correct patch identification settings, and more representation-classifier combinations.
@@ 5.4. Implication and Guidelines
We summarize the following crucial practical guidelines for future patch correctness assessment studies.
@@
Code representation is closely tied to the utilized model. The pipeline of code representation-based APCA can be typically divided into two parts: code representation for feature extraction and the corresponding model for binary prediction. The key research question lies in how to appropriately represent patched code snippets and determine the model architecture that can effectively learn the distribution of overfitting patches over correct patches. As discussed in Section 4.1, we find that one representation may vary in performance with different models and no single representation consistently achieves the optimal results, such as the accuracy of CFG ranging from 75.59% to 83.15% with three GNN models. In the future, when designing code representations, researchers are required to consider how to select more suitable models for effective source code representing and patch correctness learning.
@@
Exploring more advanced representation learning is promising. As discussed in Section 4.2, four types of code representation achieve better or comparable performance against previous state-of-the-art APCA approaches, such as an accuracy of 83.73% with the graph-based representation. Our work serves as the first attempt to empirically explore the benefits of code representation in the APCA community with some basic representations and models, such as TF-IDF and LSTM. Despite that, the preliminary results demonstrate the impressive performance in learning code semantics and predicting patch correctness. Thus, we recommend that future work be conducted in the following two directions. First, on top of existing foundational representations, it is crucial to design specific patch-oriented representation-based approaches that consider the characteristics of patches. Second, considering that a mass of code-related tasks has benefited from recent large language models (LLMs) (Tian et al., 2020; Zhang et al., 2024a), such as CodeBERT (Feng et al., 2020), it is promising to utilize the powerful code-understanding capabilities of LLMs to reason about patch correctness instead of the basic models in our work.
@@
Combining different types of code representation is challenging. As discussed in Section 4.3, it is promising to combine two types of representations, such as an accuracy improvement of 13.58% when integrating the sequence-based representation into the heuristic-based representation. However, it is challenging to combine more types of representations in our preliminary experiments, such as an only accuracy of 82.21% for the fusion of the heuristic-based, sequence-based, and tree-based representations. Future research efforts can be leveraged to explore how to better integrate a multitude of representations from two aspects. The first approach involves designing advanced fusion strategies to combine representation vectors, such as addition, multiplication, and random interleaving, while the second approach focuses on employing more powerful fusion models, such as emerging multi-modal models.
@@ 6.1. Automated Program Repair
Existing APR approaches can be categorized into four main categories according to the hypotheses to generate patches. Heuristic-based APR generates a large number of patch candidates based on syntactic program modifications and employs heuristics to explore the search space of patch generation (Le Goues et al., 2012; Martinez and Monperrus, 2016). For example, GenProg (Le Goues et al., 2012) uses genetic programming to search for a patch that passes all provided test cases and SimFix (Jiang et al., 2018) searches for code snippets from code change operations from existing patches and similar code snippets within the buggy project. Constraint-based APR transforms the patch generation into a constraint-solving problem and uses a solver to obtain a feasible solution. For example, ACS (Xiong et al., 2017) attempts to generate precise patches by ranking the ingredients for condition synthesis and is considered one of the representative approaches. Template-based APR generates patches by employing pre-defined fix patterns to mutate buggy code snippets with the retrieved donor code (Koyuncu et al., 2020; Liu et al., 2019b, a). For example, TBar (Liu et al., 2019b) summarizes a variety of fix patterns from the literature and GAMMA (Zhang et al., 2023c) is a successor of TBar by retrieving the donor code with mask prediction on top of large language models. Learning-based APR (Tufano et al., 2019; Lutellier et al., 2020; Jiang et al., 2021) treats the APR problem as a neural machine translation task with the advance of DL models. For example, Recoder (Zhu et al., 2021) predicts edits to generate syntactically-correct patches with a syntax-guided edit decoder and CIRCLE (Yuan et al., 2022) can fix multilingual bugs with a single repair model based on continual learning and T5.
@@ 6.2. Automated Patch Correctness Assessment
Learning-based APCA. Such techniques attempt to predict whether a plausible patch is correct or not based on machine learning techniques. For example, ODS (Ye et al., 2022) first extracts 202 code features at the abstract syntax tree level and trains the classifier to predict which ones are correct patches. Tian et al. (Tian et al., 2020) propose to utilize representation learning techniques to generate embedding for buggy and fix code snippets, and then input them into supervised learning classifiers to predict patch correctness. Besides, CACHE (Lin et al., 2022) is a context-aware APCA technique by considering both code context and AST structural information. Le et al. (Le-Cong et al., 2023) propose INVALIDATOR by utilizing program invariants and pre-trained models to perform semantic and syntactic reasoning for patch correctness. Recently, Zhang et al. (Zhang et al., 2024a) propose APPT, a pre-trained model-based APCA approach by jointly fine-tuning BERT and classifiers.
@@ 6.3. Code Representation in SE
The SE community has seen some studies to summarize and investigate code representation in various code-related tasks. For example, Siow et al. (Siow et al., 2022) conduct an empirical study to evaluate four categories of code representation techniques across three tasks, including code classification, vulnerability detection, and clone detection. Zhang et al. (Zhang et al., 2023a) summarize the existing learning-based APR studies into three categories according to code representation, including sequence-based, tree-based, and graph-based approaches. Similarly, Namavar (Namavar et al., 2022) invesgiate the impact of different code representations in learning-based APR approaches. Utkin et al. (Utkin et al., 2022) conduct a preliminary empirical study to explore the impact of different AST parsers on the performance of Code2Seq and TreeLSTM on the task of method name prediction. Sun et al. (Sun et al., 2023) conduct a comprehensive empirical study to explore the effectiveness of AST representation on three tasks, including clone detection, code search, and code summarization. Recently, Silva (Silva et al., 2024) et al. propose RepairLLaMA by combining realistic repair-specific code representations with parameter-efficient fine-tuning. Unlike the above studies, we are the first to explore the performance of various code representation approaches in predicting patch correctness.
@@ 7. Conclusion and Future Work
In this paper, we perform the first exploration to empirically investigate the feasibility of different code representations on automated patch correctness assessment. Specifically, we conduct a large-scale study to analyze the effectiveness and limitations of patch prediction approaches supported by code representation, involving 15 features from four categories, 11 classifiers, and more than 500 trained models. The results demonstrate that the graph-based representation can consistently outperform other representations, e.g., an accuracy of 83.95% for DDG with GGNN. We further demonstrate representation approaches achieve comparable or better performance against state-of-the-art approaches, such as CAHCE. We also explore the potential of fusing different types of code representation and discuss the impact of different node embeddings, such as textual and type node information. Lastly, our findings reveal practical guidelines for future representation-based patch correctness assessment studies. Overall, our work demonstrates the promising future and challenges of using code representation in reasoning about patch correctness and helping developers effectively exploit off-the-shelf APR tools in practice.
In the future, we will consider more powerful models and design code representations tailored for predicting patch correctness. We also hope to introduce advanced strategies to fusion different types of code representations.

### 05 Apr 2026 v2

@@ 5.1. Impact of Node Embedding
Objective. As discussed in Section 4.1, tree-based and graph-based representations have demonstrated impressive performance compared with other representations and state-of-the-art APCA approaches. The tree-based and graph-based representations consist of nodes representing code elements and edges representing relationships between nodes. In particular, each node in the tree and graph is composed of two parts, the node type and the node textual information. The former attempts to distinguish the node from the others, and the latter is a small fragment of code from the original program (the first line in each node). In our experiment, we utilize both the node type and the textual information to obtain an embedding that represents each node by default. However, it is unclear whether they both have a positive effect on the prediction results.
@@ 5.2. Potential of Pre-trained Code Models
Objective. Recently, pre-trained code models (PCMs) have achieved impressive results in software engineering tasks, such as program repair and test generation (Zhang et al., 2026a). Since the strong performance of these models may come not only from code representations but also from large-scale pre-training and stronger model capacity, we did not include them in the main experiments, whose goal is to study the impact of code representations under a more controlled setting. In this section, we further examine the potential of PCMs for predicting patch correctness.
Design. We select two representative PCMs, i.e., the encoder-only CodeBERT and the encoder-decoder CodeT5. As both models are pre-trained on code sequences, we represent the input code as token sequences and feed them into the models. For comparison, we also include Transformer, the best-performing non-pre-trained sequence-based model in our study.
Results. Table 9 shows the results. Overall, both PCMs achieve competitive performance, indicating their potential to learn patch semantics and assess patch correctness. Among the three models, CodeBERT achieves the best overall performance, with 83.22% accuracy, 82.48% precision, 83.48% recall, 82.93% F1, and 87.80% AUC. Compared with Transformer, CodeBERT improves accuracy, precision, F1, and AUC by 0.74%, 3.51%, 0.06%, and 5.20%, respectively, while yielding a lower recall. CodeT5 also performs competitively, achieving an average score of 83.61%, which is 0.77% higher than Transformer. These results suggest that PCMs can be effective for APCA and may offer stronger discriminative ability than non-pre-trained sequence models. However, the improvements are still modest. More importantly, PCM-based and representation-oriented approaches are largely orthogonal: the former emphasizes stronger model architectures and pre-training, whereas the latter focuses on better encoding patch semantics. Therefore, an important direction for future work is to combine advanced PLMs with more effective code representations to assess patch correctness.
@@ 5.3. Threats to Validity
The first threat to validity is the selection of baselines. In Section 4.2, we consider four APCA techniques as baselines to evaluate the effectiveness of four types of representations. We may fail to consider other techniques (mentioned in Section 6) due to the rapidly growing community and the constant emergence of APCA approaches. However, our selected baselines represent state-of-the-art (e.g., the traditional dynamic-based PATCH-SIM) and are widely-adopted in almost all of previous studies (Lin et al., 2022; Tian et al., 2020). Besides, these selected baselines cover various categories, such as ODS (Ye et al., 2022) with hand-crafted features, Tian et al. (Tian et al., 2020) with code sequences, and CACHE (Lin et al., 2022) with ASTs. Furthermore, our work mainly focuses on empirical evaluations of the effectiveness of different representations. Thus, we believe the baselines have little impact on our results.
@@
The second threat to validity is the evaluation benchmark. In our experiment, we focus on the most popular Defects4J in the APR and APCA literature involving 835 bugs from 17 open-source projects. It is unclear to what degree the results in our work can be generalized to other benchmarks. Despite other datasets available (Lin et al., 2022), they are quite easy to identify due to simple bug types, such as 99.9% AUC (Zhang et al., 2024a), making it difficult to assess true capabilities. To address this threat, we collect 2,274 patches generated by 39 existing APR tools, to the best of our knowledge, which is the largest patch set of Defects4J in the literature. It should be noted that, with the continuous release of newer APR tools on Defects4J, we observed that their contributions of unique patches (i.e., those not already generated by earlier tools) have become increasingly limited. In fact, some of their patches overlap with patches already included in our dataset. For this reason, while Table 1 does not explicitly list these tools, their outcomes are partially covered, and our dataset remains both comprehensive and the largest available. Besides, we mitigate the potential bias by using multiple evaluation metrics to exhaustively assess the involved techniques. In the future, we plan to continuously update the dataset by incorporating newer APR tools, such as Tare (Zhu et al., 2023) and Recoder (Zhu et al., 2021), to further strengthen the completeness of our benchmark.
@@
The third threat to validity lies in the evaluation setting. First, following most previous studies (Lin et al., 2022; Tian et al., 2020), we evaluate the performance of APCA approaches within a cross-validation setting, which is the common practice in the APCA community. It is possible that patches generated by different APR tools for the same bug are split between the training and testing sets. However, in a real-world scenario, it would be impossible to use the label from one patch to predict the other for the same bug. Second, there exist more potential code representations and models that are excluded in our experiments due to page limit. Third, we focus on filtering overfitting patches following most prior work and the results may not be generalized to correct patch identification. To address this threat, we construct a large benchmark with 2,274 patches from 39 APR tools, consider 15 representations and 11 models, and select seven baselines from different categories. Besides, considering the scale of our empirical evaluation involving nearly 600 trained models, i.e., (45 models in RQ1 + one model in RQ2 + 22 models in RQ3 + 51 models in Section 5.1) $\times$ 5-fold cross-validation, the current results are enough to demonstrate the performance of code representations in patch correctness prediction. In the future, researchers could further explore the experiments under more potential scenarios, such as the cross-project evaluation and correct patch identification settings, and more representation-classifier combinations.
@@ 5.4. Implication and Guidelines
We summarize the following practical guidelines for future studies on patch correctness assessment.
@@
Code representation is closely tied to the utilized model. The pipeline of code representation-based APCA can typically be divided into two parts: code representation for feature extraction and the corresponding model for binary prediction. The key research question lies in how to appropriately represent patched code snippets and determine the model architecture that can effectively learn the distribution of overfitting patches over correct patches. As discussed in Section 4.1, we find that one representation may vary in performance with different models and no single representation consistently achieves the optimal results, such as the accuracy of CFG ranging from 75.59% to 83.15% with three GNN models. In the future, when designing code representations, researchers are required to consider how to select more suitable models for effective source code representation and patch correctness learning.
@@
Exploring more advanced representation learning is promising. As discussed in Section 4.2, four types of code representation achieve better or comparable performance against previous state-of-the-art APCA approaches, such as an accuracy of 83.73% with the graph-based representation. Our work serves as the first attempt to empirically explore the benefits of code representation in the APCA community with some basic representations and models, such as TF-IDF and LSTM. Despite that, the preliminary results demonstrate the impressive performance in learning code semantics and predicting patch correctness. Thus, we recommend that future work be conducted in the following two directions. First, on top of existing foundational representations, it is crucial to design specific patch-oriented representation-based approaches that consider the characteristics of patches. Second, given that a large body of code-related tasks has benefited from recent large language models (Tian et al., 2020; Zhang et al., 2024a), it is promising to leverage LLMs’ powerful code-understanding capabilities to reason about patch correctness, rather than relying on basic models in our work.
@@
Combining different types of code representation is challenging. As discussed in Section 4.3, it is promising to combine two types of representations, such as an accuracy improvement of 13.58% when integrating the sequence-based representation into the heuristic-based representation. However, it is challenging to combine more types of representations in our preliminary experiments, such as an accuracy of only 82.21% for the fusion of the heuristic-based, sequence-based, and tree-based representations. Future research efforts can be leveraged to explore how to better integrate a multitude of representations from two aspects. The first approach involves designing advanced fusion strategies to combine representation vectors, such as addition, multiplication, and random interleaving, while the second approach focuses on employing more powerful fusion models, such as emerging multi-modal models.
@@ 6.1. Automated Program Repair
Existing APR approaches can be categorized into four main categories according to the hypotheses to generate patches. Heuristic-based APR generates a large number of patch candidates based on syntactic program modifications and employs heuristics to explore the search space of patch generation (Le Goues et al., 2012; Martinez and Monperrus, 2016). For example, GenProg (Le Goues et al., 2012) uses genetic programming to search for a patch that passes all provided test cases and SimFix (Jiang et al., 2018) searches for code snippets from code change operations from existing patches and similar code snippets within the buggy project. Constraint-based APR transforms the patch generation into a constraint-solving problem and uses a solver to obtain a feasible solution. For example, ACS (Xiong et al., 2017) generates precise patches by ranking the ingredients for condition synthesis and is considered one of the representative approaches. Template-based APR generates patches by employing pre-defined fix patterns to mutate buggy code snippets with the retrieved donor code (Koyuncu et al., 2020; Liu et al., 2019b, a). For example, TBar (Liu et al., 2019b) summarizes a variety of fix patterns from the literature and GAMMA (Zhang et al., 2023b) is a successor of TBar by retrieving the donor code with mask prediction on top of LLMs. Learning-based APR (Tufano et al., 2019; Lutellier et al., 2020; Jiang et al., 2021) treats the APR problem as a neural machine translation task with the advance of DL models. For example, Recoder (Zhu et al., 2021) predicts edits to generate syntactically-correct patches with a syntax-guided edit decoder and CIRCLE (Yuan et al., 2022) can fix multilingual bugs with a single repair model based on continual learning and T5. Furthermore, recent studies have explored the use of LLMs for program repair (Xia et al., 2025; Zhang et al., 2026b; Applis et al., 2026), which can be viewed as a further extension of learning-based APR. Since this direction is not the main focus of this paper, we do not discuss it in detail; a comprehensive overview is available in the survey (Zhang et al., 2026a).
@@ 6.2. Automated Patch Correctness Assessment
Learning-based APCA. Such techniques attempt to predict whether a plausible patch is correct or not based on machine learning techniques (Zhou et al., 2024). For example, ODS (Ye et al., 2022) first extracts 202 code features at the abstract syntax tree level and trains the classifier to predict which ones are correct patches. Tian et al. (Tian et al., 2020) propose to utilize representation learning techniques to generate an embedding for buggy and fixed code snippets, and then input them into supervised learning classifiers to predict patch correctness. Besides, CACHE (Lin et al., 2022) is a context-aware APCA technique by considering both code context and AST structural information. Le et al. (Le-Cong et al., 2023) propose INVALIDATOR by utilizing program invariants and pre-trained models to perform semantic and syntactic reasoning for patch correctness. Recently, Zhang et al. (Zhang et al., 2024a) propose APPT, a pre-trained model-based APCA approach by jointly fine-tuning BERT and classifiers.
@@ 6.3. Code Representation in SE
The SE community has seen some studies to summarize and investigate code representation in various code-related tasks. For example, Siow et al. (Siow et al., 2022) conduct an empirical study to evaluate four categories of code representation techniques across three tasks, including code classification, vulnerability detection, and clone detection. Zhang et al. (Zhang et al., 2023a) summarize the existing learning-based APR studies into three categories according to code representation, including sequence-based, tree-based, and graph-based approaches. Similarly, Namavar (Namavar et al., 2022) invesgiate the impact of different code representations in learning-based APR approaches. Utkin et al. (Utkin et al., 2022) conduct a preliminary empirical study to explore the impact of different AST parsers on the performance of Code2Seq and TreeLSTM on the task of method name prediction. Sun et al. (Sun et al., 2026) conduct a comprehensive empirical study to explore the effectiveness of AST representation on three tasks, including clone detection, code search, and code summarization. Recently, Silva (Silva et al., 2024) et al. propose RepairLLaMA by combining realistic repair-specific code representations with parameter-efficient fine-tuning. Unlike the above studies, we are the first to explore the performance of various code representation approaches in predicting patch correctness.
@@ 7. Conclusion and Future Work
In this paper, we perform the first exploration to empirically investigate the feasibility of different code representations on automated patch correctness assessment. Specifically, we conduct a large-scale study to analyze the effectiveness and limitations of patch prediction approaches supported by code representation, involving 15 features from four categories, 11 classifiers, and more than 500 trained models. The results demonstrate that the graph-based representation can consistently outperform other representations, e.g., an accuracy of 83.95% for DDG with GGNN. We further demonstrate representation approaches achieve comparable or better performance against state-of-the-art approaches, such as CAHCE. We also explore the potential of fusing different types of code representation and discuss the impact of different node embeddings, such as textual and type node information. Lastly, our findings reveal practical guidelines for future representation-based patch correctness assessment studies. Overall, our work demonstrates the promising future and challenges of using code representation in reasoning about patch correctness and helping developers effectively exploit off-the-shelf APR tools in practice. In the future, we will consider more powerful models and design code representations tailored for predicting patch correctness. We also hope to introduce advanced strategies to fuse different types of code representations.